# Quasi-randomization tests for network interference

### v2

@@ Appendix A Algorithm: $\epsilon$-net cluster
end while
Store the resultant cluster outputs in $C$
Permute the outputs to the original ordering and store them in $C,V$
Output: C and V
@@ Appendix B Proofs
B.1 Proof of Theorem 1:
@@
$p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs}).$
Given $Z_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution induced by $P(G|Z_{obs})$. Thus,
$p_{Z_{obs}}=1-F_{U}(G_{obs}|Z_{obs}).$
@@
Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z_{obs})$ and we get
@@
B.2 Proof of Theorem 2 and Theorem 3:
@@
Given $Z_{obs}$ and $[G_{obs}]$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution as induced by $P(G|Z=Z_{obs},G\in[G_{obs}])$. Thus,
@@
Let $P_{1},P_{2},...,P_{d}$ be the partition induced by $\equiv_{d}$. Pick $G_{1},G_{2},...,G_{d}$ such that $G_{i}\in P_{i}$ for all $i=1,2,...,d$. Hence,
@@
Let $[G_{obs}]=\{G\in\{0,1\}^{N\times N}:\>deg(G)=deg(G_{obs})\}$. We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}])$ is a uniform distribution under a completely randomized experiment. Consider $g,g^{\prime}\in[G_{obs}]$.
@@
$[G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\>|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
Here, ${Z_{obs}}^{t}=\{i\in\mathbb{P}:\>{Z_{obs}}_{i}=t\}$ for $t\in\{0,1\}$. Note that $[G_{obs}|Z_{obs}]\subseteq[G_{obs}]$, where $[G_{obs}]=\{G\in\mathcal{G}|G\cong G_{obs}\}$. We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}|Z_{obs}])$ is a uniform distribution under a cluster randomized experiment. Let $C$ be a random variable defined as $\mathcal{A}\circ G$. Here, $\mathcal{A}$ represents the clustering algorithm in Algorithm 1.
@@
$=P(\sum_{i=1}^{|C|}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\;j\in[N]\>|\>G=g,G\in[G_{obs}|Z_{obs}])$
$=P(\sum_{i=1}^{|\pi(C_{obs})|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))={Z_{obs}}_{j}\;j\in[N]\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(\pi^{-1}(j)\in{C_{obs}}_{i})=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in{C_{obs}}_{i})=t\quad\forall\pi(j)\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
$=P(W=W_{obs})=P(Z=Z_{obs}|G=g^{\prime},G\in[G_{obs}|Z_{obs}]).$
@@ C.1 Erdős-Rényi random graph model
We formally describe the Erdős-Rényi random graph model (Erdős et al. [1960]). The below-described model is the probabilist variant of the (Erdős-Rényi) random graph model. We consider the number of nodes, $N$, and edges, $|E|$, to be fixed, and any two nodes will be linked independently with an edge with a fixed probability $p$. That is,
@@
To model $P(G)$ using Equation (14), we have to estimate the parameter $p$. We obtain the MLE of the parameter $p$ and estimate it as $\frac{|E|}{\binom{N}{2}}$. We conduct an extensive numerical simulation study and verify the finite sample validity of the proposed randomization test. We also obtain the power of the test under different scenarios. An Erdős-Rényi random graph model is the simplest network model one can consider, and it may not be representative of the kind of networks we observe in the real world. For example, one observes a high degree of clustering in social networks in the graph (Watts and Strogatz [1998]). One drawback of the model is the independence and identical probability of formation of two edges in the graph. One way to characterize this is to evaluate the degree distribution of the random graph. It has been shown that the degree distribution of a large Erdős-Rényi random graph is a Poisson distribution. Most real-life networks violate this property and have highly skewed degree distribution. In the main text, we generalize this and characterize random graphs with arbitrary degree distribution.
@@ C.2 Data generating mechanism
$Y_{i}(\textbf{z})=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot z_{j}>0)+\epsilon_{i},$ (15)
@@
We generate the randomization distribution of the test statistic in consideration by repeated drawing from the Erdős-Rényi random graph model with the probability of an edge between two nodes being 0.2. The simulations in Table 3 were carried out on 100 units, with 5 being treated and the rest in control. We approximated the randomization distribution of the test statistic by taking 200 samples of the graph. We calculate the corresponding p-value by the proportion of times the test statistic imputed was larger than the test statistic for the observed network. If the p-value is less than 0.05, we reject the null hypothesis of no spillover effects. We replicated the above 450 times and calculated rejection rates as the proportion of times the p-value was significant. For this setup, the outcome model (15) with $\tau_{direct}$ equals to 4 and $\tau_{spill}$ is presented in Table 3, varying from $0.0$ to $0.7$.
We then repeat the simulations for the above setup for outcome model (16) and present them in Table 4. We note that the power of the test decreases substantially for model (16), compared with model (15). This is owing to the simplistic nature of the outcome model (15), and we present its results for verification purposes. Another caveat in the above setting is the nature of test statistic $T_{I_{c}}(G|Z_{obs})$. We might not get any control units with all neighboring ones in control depending on the input graph, invalidating the procedure. So, the above setup involves 100 units with 5 units in treatment. We numerically found the procedure failing when the number of treated units was increased to more than 5. We present the results for model (16) with $T_{quant}$ as test statistic in Table 4. Here, the number of units in the population is 500, with the number of treated units being 250. $\tau_{direct}$ is 4 and varying $\tau_{spill}$ is used for power performance measure.
In all the above models, we took the probability of an edge between two nodes as $0.2$. We present the same setup in 4, with the probability of an edge between two nodes estimated below. $G_{obs}$ is generated from Erdős-Rényi random graph with $p=0.2$.

### v3

@@ Appendix A Algorithm: $\epsilon$-net cluster
  • Step 4: Store the resultant cluster outputs in $C$
  • Step 5: Permute the outputs to the original ordering and store them in $C,V$
  • Output C and V
@@ Appendix B Proofs
    B.1 Proof of Proposition 1
    Proof.
    Note that under $\mathcal{H}_{0}$,
    $Y_{{G_{X}}_{i}}(Z,G)=Y_{{G^{\prime}_{X}}_{i}}(Z^{\prime},G^{\prime})\qquad Z_{i}=Z^{\prime}_{i},{G_{X}}_{i}={G^{\prime}_{X}}_{i}.$
    Thus,
    $T_{c}(G|Z_{obs})$ $=T(Z_{obs},G,Y_{{G_{X}}_{obs}}(Z_{obs},G)),$
    $=T(Z_{obs},G,Y_{{G_{X}}_{obs}}(Z_{obs},G^{\prime}))\qquad P(G^{\prime}|Z_{obs},{G_{X}}_{obs})>0\text{ under }\mathcal{H}_{0}.$
    Since $Z_{obs}$ is arbitrary, we conclude imputability of the $T_{c}(G|Z_{obs})$ under the null $\mathcal{H}_{0}$. Take $G^{\prime}=G_{obs}$, we get $T_{c}(G|Z_{obs})=T(Z_{obs},G,Y_{obs})$. ∎
    B.2 Proof of Proposition 2
    Proof.
    B.3 Proof of Theorem 1
@@
    $p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs},{G_{X}}_{obs}).$
    Given $Z_{obs}$ and ${G_{X}}_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution induced by $P(G|Z_{obs},{G_{X}}_{obs})$. Thus,
    $p_{Z_{obs}}=1-F_{U}(G_{obs}|Z_{obs},{G_{X}}_{obs}).$
@@
    Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs},{G_{X}}_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z_{obs},{G_{X}}_{obs})$ and we get
@@
    B.4 Proofs of Theorem 2 and Theorem 3
@@
    Given $Z_{obs}$ and $G_{obs}]$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution as induced by $P(G|Z=Z_{obs},G\in[G_{obs}])$. Thus,
@@
    Let $P_{1},P_{2},...,P_{d}$ be the partition induced by $\equiv_{d}$. Pick $G_{1},G_{2},...,G_{d}$ such that $G_{i}\in P_{i}$ for all i=1,2,,di=1,2,...,d. Hence,
@@
    Let $G_{obs}]=\{G\in\{0,1\}^{N\times N}:\>deg(G)=deg(G_{obs})\>s.t.\>deg_{i}(G)=deg_{i}(G_{obs})\;\forall i\in[N]\}$. We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}])$ is a uniform distribution under a completely randomized experiment. Consider $g,g^{\prime}\in[G_{obs}]$.
@@
    $G_{obs}|Z_{obs}]=\{G\in[G_{obs}]\>|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
    Here, ${Z_{obs}}^{t}=\{i\in\mathbb{P}:\>{Z_{obs}}_{i}=t\}$ for $t\in\{0,1\}$. Note that $G_{obs}|Z_{obs}]\subseteq[G_{obs}]$, where $G_{obs}]=\{G\in\mathcal{G}|G\cong G_{obs}\}$. We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}|Z_{obs}])$ is a uniform distribution under a cluster randomized experiment. Let $C$ be a random variable defined as $\mathcal{A}\circ G$. Here, $\mathcal{A}$ represents the clustering algorithm in Algorithm 1.
@@
    $\hskip-85.35826pt=P(\sum_{i=1}^{|C|}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\;j\in[N]\>|\>G=g,G\in[G_{obs}|Z_{obs}])$
    $\hskip-85.35826pt=P(\sum_{i=1}^{|\pi(C_{obs})|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))={Z_{obs}}_{j}\;j\in[N]\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
    $\hskip-85.35826pt=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
    $\hskip-85.35826pt=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(\pi^{-1}(j)\in{C_{obs}}_{i})=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
    $\hskip-85.35826pt=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in{C_{obs}}_{i})=t\quad\forall\pi(j)\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
    $\hskip-85.35826pt=P(W=W_{obs})=P(Z=Z_{obs}|G=g^{\prime},G\in[G_{obs}|Z_{obs}]).$
@@ C.1 Erdős-Rényi random graph model
    We formally describe the Erdős-Rényi random graph model (Erdős et al. [1960]). The model described below is the probabilist variant of the (Erdős-Rényi) random graph model. We consider the number of nodes, $N$, and edges, $|E|$, to be fixed, and any two nodes will be linked independently with an edge with a fixed probability $p$. That is,
@@
    To model $P(G)$ using Equation (15), we have to estimate the parameter $p$. We obtain the MLE of the parameter $p$ and estimate it as $\frac{|E|}{\binom{N}{2}}$. We conduct an extensive numerical simulation study and verify the finite sample validity of the proposed randomization test. We also obtain the power of the test under different scenarios. An Erdős-Rényi random graph model is the simplest network model one can consider, and it may not be representative of the kind of networks we observe in the real world. For example, one observes a high degree of clustering in social networks in the graph (Watts and Strogatz [1998]). One drawback of the model is the independence and identical probability of formation of two edges in the graph. One way to characterize this is to evaluate the degree distribution of the random graph. It has been shown that the degree distribution of a large Erdős-Rényi random graph is a Poisson distribution. Most real-life networks violate this property and have highly skewed degree distributions. In the main text, we generalize this and characterize random graphs with an arbitrary degree distribution.
@@ C.2 Data generating mechanism
    $Y_{i}(z)=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot z_{j}>0)+\epsilon_{i},$ (16)
@@
    We generate the randomization distribution of the test statistic under consideration by repeated drawing from the Erdős-Rényi random graph model with the probability of an edge between two nodes being 0.2. The simulations in Table 4 were carried out on 100 units, with 5 being treated and the rest in control. We approximated the randomization distribution of the test statistic by taking 200 samples of the graph. We calculate the corresponding p-value by the proportion of times the test statistic imputed was larger than the test statistic for the observed network. If the p-value is less than 0.05, we reject the null hypothesis of no spillover effects. We replicated the above 450 times and calculated rejection rates as the proportion of times the p-value was significant. For this setup, the outcome model (16) with $\tau_{direct}$ equals to 4 and $\tau_{spill}$ is presented in Table 4, varying from $0.0$ to $0.7$.
    We then repeat the simulations for the above setup for outcome model (17) and present them in Table 5. We note that the power of the test decreases substantially for model (17), compared with model (16). This is owing to the simplistic nature of the outcome model (16), and we present its results for verification purposes. Another caveat in the above setting is the nature of the test statistic $T_{I_{c}}(G|Z_{obs})$. We might not get any control units with all neighboring ones in control, depending on the input graph, invalidating the procedure. So, the above setup involves 100 units with 5 units in treatment. We numerically found the procedure failing when the number of treated units was increased to more than 5. We present the results for model (17) with $T_{quant}$ as test statistic in Table 5. Here, the number of units in the population is 500, with the number of treated units being 250. $\tau_{direct}$ is 4 and varying $\tau_{spill}$ is used for power performance measure.
    In all the above models, we took the probability of an edge between two nodes as $0.2$. We present the same setup in 5, with the probability of an edge between two nodes estimated below. $G_{obs}$ is generated from Erdős-Rényi random graph with $p=0.2$.